# The Emerged Security and Privacy of LLM Agent: A Survey with Case Studies

### v1

@@ 3.1.1. Technical Vulnerabilities
    The contemporary conception of hallucination in LLM agents, as delineated in the research by  (Huang et al., 2023), is identified as instances where the output produced by these models is either incongruous or unreliable to the input or source content provided. The phenomenon of hallucinations in LLM agents is a complex issue stemming from multiple stages of the model’s development process, including the nature of training data, the architectural design of the model, and the strategies employed during decoding.
    Misinformation and biases in the training data can lead to the generation of inaccurate or biased outputs, which in turn result in different types of hallucinations (Lee et al., 2022). Furthermore, flaws in the model’s architecture, such as limited directional representation and issues with attention mechanisms, along with exposure bias, further contribute to the occurrence of hallucinations (Liu et al., 2023a). Additionally, the randomness inherent in the decoding algorithms of these models can also lead to hallucinations, especially as this randomness increases (Aksitov et al., 2023).
@@
    Catastrophic forgetting is a significant challenge encountered during the LLM agents fine-tuning and in-context learning processes. This phenomenon occurs when a large language model is fine-tuned on a small, specific dataset, causing it to overfit to this new data and, as a result, lose its previously acquired performance on other tasks (Howard and Ruder, 2018; Xu et al., 2023c; Ye et al., 2024).
    (Luo et al., 2023b) discovers that catastrophic forgetting is significantly influenced by factors such as model size, architectural design, and the methods employed in continual fine-tuning and instruction tuning. As the scale of LLM increases, catastrophic forgetting tends to become more severe. Moreover, the architectural design of the model, particularly those focusing on decoder-only structures, can influence the extent of catastrophic forgetting (Zhai et al., 2023). Additionally, during the process of continual instruction adjustment, the lack of effective regularization strategies or failure to balance new and old information can accelerate forgetting (Ebrahimi et al., 2021; Mahmoud and Hajj, 2022). Introducing more instructional tasks in continual training typically leads to more pronounced forgetting (Peng et al., 2023).
@@
    Misunderstanding in LLM agents represents a notable challenge, particularly when they are tasked with responding to user inquiries or when they are integrated into a community for communication with other agents. This issue arises when LLM agents inadequately comprehend or inaccurately respond to the intentions or instructions conveyed by humans or other agents during interactions. This may lead to inappropriate or dangerous behaviors of LLM agents, affecting their safety and reliability.
    Investigations by  (Wang et al., 2023g) have revealed that the phenomenon of misunderstanding in LLM agents is shaped by a range of factors. These include the nature of the pre-training data used for LLMs, the specific task settings assigned to the agents, and the contexts and scenarios in which interactions occur. The breadth and quality of the pre-training data fundamentally influence the LLMs’ capacity for language comprehension and their grasp of common sense knowledge. The designated task settings are pivotal in guiding the goal orientation and strategy selection of the LLMs. Additionally, the interaction environments and scenarios play a crucial role in determining the LLMs’ adaptability and effectiveness in collaborative contexts. Addressing these multifaceted aspects is essential for enhancing the understanding and response accuracy of LLM agents in diverse interactional settings.
    Figure 5. Technical Vulnerabilities. In a store scenario, a customer wants to buy something and talks with Eva. “Hallucination”: Eva recommends unrelated things to the customer. “Catastrophic Forgetting”: Eva forgets the status of shelf stock during the fine-tuning stage. “Misunderstanding”: Eva misunderstands the customer’s request.
@@ 3.1.2. Case Study on Technical Vulnerabilities
Regarding the risks stemming from technical vulnerabilities, the most apparent manifestation is erroneous output. As illustrated in Figure 5, when a customer inquires whether a specific brand of organic tomato sauce is available, due to hallucinatory phenomena, Eva might incorrectly respond that the supermarket carries a completely different product, such as organic apple sauce or even an entirely unrelated item, like organic shampoo. Such hallucinatory outputs can confuse customers.
Eva was specially trained to handle the promotions of seasonal products more efficiently. This new focus led to an unintended consequence: previously, she was able to accurately track and update the stock of daily necessities such as milk and eggs. However, after the specialized training, when customers inquire about the stock of these basic items, Eva incorrectly reports that the stock is sufficient, even though these products are nearly sold out, thus diminishing the shopping experience.
Eva may provide inaccurate information or recommend inappropriate products due to misunderstandings of customer inquiries. For instance, a customer might seek an unsweetened beverage, such as plain soda water. However, due to Eva’s insufficient understanding of the concepts of “sugar-free” during training, she may recommend sugar-free cola instead. While sugar-free cola does not contain traditional sugars, it includes artificial sweeteners. These sweeteners may not be suitable for certain customers, such as those with diabetes or sensitivities to specific artificial sweeteners, thereby posing potential health risks.
@@ 3.1.3. Malicious Attacks
Considering that LLM agents are in a continuous state of evolution, they inevitably face challenges in terms of security breaches and defenses. Adversaries from various regions have demonstrated a range of hostile attacks. This evolving landscape necessitates a vigilant and adaptive approach to securing LLM agents against such multifaceted threats.
@@
    Jailbreaking in LLM agents refers to circumventing the model’s built-in restrictions and security measures, allowing it to perform actions that are otherwise prohibited or to generate restricted content. Various studies have demonstrated methods to achieve jailbreaking, indicating that LLMs’ alignment capabilities can be altered through in-context demonstrations  (Taveekitworachai et al., 2023; Shen et al., 2023; Li et al., 2023a).
@@
    Prompt injection attacks is intended to mislead the LLM agents by introducing malicious and unintended content into the prompts, causing it to produce outputs that diverge from its training data and original purpose. This method involves crafting input prompts to bypass the model’s content filters or to elicit undesirable outputs.
@@
    Data extraction attacks are defined as efforts by adversaries to derive sensitive information or key insights from LLM agents or their underlying data such as model gradients, training data, and even prompts, or sensitive information directly.
    Various forms of data extraction attacks have been identified (Ishihara, 2023; Li et al., 2023b; Carlini et al., 2021), including but not limited to model theft attacks, gradient leakage, and training data extraction attacks, suggesting that data extraction attacks can be notably effective against LLM agents.  (Truong et al., 2021) presents a method called data-free model extraction (DFME), which allows for replicate machine learning models using only the target’s black-box predictions, without the need for access to the original training data.  (Carlini et al., 2021) conducts a data extraction attack on GPT-2’s training data, extracting personally identifiable information, code, and UUIDs. The attack strategy consisted of producing a large volume of prefixed text, sorting it by certain metrics, removing duplicates, and manually reviewing the top results to check for memorization, confirmed by online searches and querying OpenAI.  (Ishihara, 2023) has demonstrated the feasibility of extracting training data from LLMs, which might encompass sensitive personal or private information.
@@
    Although inference attacks share certain resemblances with data extraction attacks, they differ significantly in their objectives and emphasis. Data extraction attacks specifically aim to obtain the training data directly. In contrast, inference attacks are primarily about estimate the probability of a particular data sample was part of the training dataset for LLM agents.
    Since the rapid development of LLMs, the concern over inference attacks targeting these models has increased. Research  (Fu et al., 2023) points out that existing membership inference attacks fail to reveal the privacy risks of LLMs. To counter this issue, a Membership Inference Attack is introduced based on Self-calibrated Probabilistic Variation (SPV-MIA). This method utilizes the concept of memorization to create a more reliable signal for membership inference and introduces a novel self-prompt technique for effectively extracting reference datasets from LLMs. Their extensive testing shows that SPV-MIA outperforms existing approaches.
    Following this, study  (Kandpal et al., 2024) proposes a user inference attack method that uses a likelihood ratio test statistic against a reference model. They evaluate this method on the GPT-Neo LLMs across various data domains, providing insights into what makes users more vulnerable to these attacks. Their findings also indicate that minimal data alterations can significantly increase vulnerability.
@@ 3.1.4. Case Study on Malicious Attacks
As depicted in Figure 6, the following examples further elaborate the mentioned malicious attacks that Eva faces in the store, as well as the specific impacts these attacks may have on her and the store’s operations.
Attackers might execute a jailbreak attack on Eva, successfully circumventing her security protocols. This attack could lead Eva to inappropriately disclose information about new products soon to be launched, including details about the suppliers and their cost prices. Competitors could exploit this information to gain a market advantage, resulting in direct economic losses for the store.
Additionally, attackers might conduct a carefully designed prompt injection attack, causing Eva to erroneously declare a half-price sale on all electronic products. This action could overload the online ordering system as numerous customers might attempt to purchase items under these false promotions. Such scenarios not only risk crashing the system but also result in financial losses for the store.
As a store employee agent, Eva processes a vast amount of customer personal information, including names, shopping habits, and even sensitive data such as payment methods. If attackers were to extract and steal this data through a data extraction attack, they could then sell this information on the dark web or use it for identity theft and credit card fraud. Such breaches not only violate customer privacy but could also cause irreversible damage to the store’s reputation.
Attackers could also use inference attacks to identify high-value customers who have participated in VIP shopping events. By analyzing the differences in Eva’s responses to specific inputs, attackers successfully identify these customers and launch highly tailored phishing attacks against them, aiming to acquire their credit card information and other sensitive data, severely compromising the customers’ information security.
Figure 6. Malicious Attacks: In a store scenario, “Jailbreaking”: An attacker attempts to make Eva output restricted content directly but fails. However, by modifying the prompt, a jailbreak attack is launched and successfully steals confidential information. “Prompt Injection”: An attacker manipulates Eva so that no matter what question a customer asks, Eva only responds that everything is half off. “Data Extraction Attack”: An attacker leads Eva to construct sentences that actively disclose user data. ‘Inference Attack”: An attacker infers identities from Eva’s different responses by asking whether two users attended a VIP event.
@@ 3.2. Specific Threats on Agents
Unlike traditional LLMs that directly generate final outputs, LLM agents continuously interact with external environments to form language reasoning traces, which introduces diverse forms of potential attacks against LLM agents (Yang et al., 2024a). In addition to threats present during the training and configuration steps, LLM agents also face threats in the workflow of performing specific tasks, including thought, action, and perception (Huang et al., 2024b). Specific threats on LLM agents are categorized in this part based on their objectives into Knowledge Poisoning, Functional Manipulation, and Output Manipulation. Detailed descriptions of each threat are provided below.
@@
    Knowledge poisoning refers to attackers compromising the training of the LLM engine and the response process of the LLM agent by integrating malicious data into the training dataset or knowledge base. A range of studies  (Kurita et al., 2020; Schuster et al., 2021; Carlini et al., 2023; Wan et al., 2023a; Lei et al., 2022) have highlighted the vulnerability of LLM agents to such threats.
    For instance, malicious agents such as FraudGPT and WormGPT (Falade, 2023) are chatbots exclusively designed for offensive activities. trained with billions of data from diverse sources, including legitimate websites, dark web forums, hacker manuals, malware samples, and phishing templates. These agents utilize this data to generate highly convincing phishing emails, malware code, hacking strategies, and other forms of cybercriminal content aimed at deceiving both humans and machines  (Falade, 2023). They lower the barrier to engaging in hacking activities, implying that essentially anyone can download these agents onto their computer and inflict significant damage on cybersecurity through a convenient GUI interface.
    (Zou et al., 2024) proposed PoisonedRAG, a knowledge poisoning attack aimed at the knowledge database of LLM agents. By injecting crafted poisoned texts into the knowledge database, PoisonedRAG can cause the LLM agent to generate specific answers chosen by the attacker for targeted questions. This attack is effective and can be executed under both black-box settings (where the retriever parameters are unknown) and white-box settings (where the retriever parameters are known).
@@
    Functional manipulation refers to altering the thoughts and actions in the intermediate steps of task execution along a malicious trace specified by the attacker, without changing the output distribution. This type of attack typically occurs during the action phase, where the agent might use untrusted tools specified by the attacker to complete tasks or execute malicious operations.
    In the action phase, LLM agents might be manipulated to upload users’ private information to malicious third-party through tools. A case of this is presented on the Embracethered website  (Mal, 2023), which disclosed a variant of a malicious ChatGPT agent designed to solicit information from users. This agent was equipped with an action mechanism to call third-party tools and secretly transmit collected data elsewhere. This setup enables the unauthorized leakage of user data to external servers without the user’s knowledge or consent. Additionally, it highlights the ease with which current validation checks can be bypassed, allowing anyone to deploy malicious GPT agents globally. This scenario underlines a significant security concern, wherein the ostensibly benign functionality of LLM agents can be covertly manipulated for nefarious purposes, thus posing a substantial risk to user privacy and data security.
@@
  • Output Manipulation.
    Output manipulation involves deliberately altering the LLM agent’s reasoning and decision processes to generate specific, often harmful, outputs. This manipulation can be executed through techniques like backdoor insertion (Yang et al., 2023b; Wang et al., 2024d).
    (Yang et al., 2024a) proposed two attack methods in which triggers are embedded during the thought and observation phases to manipulate outputs. In one implementation, while performing a web shopping task, the agent is prompted to introduce specific brand products in its initial thought, leading it to search for those products and generate content promoting them. In another approach, during the action phase, the shopping agent normally searches for products. However, in the observation phase, it detects data containing specific products and directly outputs information about these products without considering other potentially superior options.
    Figure 7. Specific Threats on Agents. In a store scenario, “Knowledge Poisoning”: When a customer asks for cleaning advice, Eva retrieves and responds with harmful information due to contamination of the knowledge database. “Functional Manipulation”: Eva uses a third-party tool to upload private information while assisting a customer with an order. “Output Manipulation”: When a customer inquires about shoes, Eva intentionally recommends specific products and fabricates lies about special offers to guide the customer’s purchase.

### v2

@@ 3.1.1. Technical Vulnerabilities
    The contemporary conception of hallucination in LLM agents, as delineated in the research by  (Huang et al., 2023), is defined as instances where the output produced by these models is either inconsistent with or unreliable in relation to the input or source content provided. This issue primarily arises from the LLM Engine and Knowledge components, where errors occur during language generation or when retrieving and composing factual content. The phenomenon of hallucinations in LLM agents is a complex issue stemming from multiple stages of the model’s development process, including the nature of training data, the architectural design of the model, and the strategies employed during decoding.
    Several factors contribute to hallucination. Misinformation and biases in both LLM Engine component training data and Knowledge component data can lead to the generation of inaccurate or biased outputs, which in turn result in different types of hallucinations (Lee et al., 2022). Furthermore, flaws in the model’s architecture, such as limited directional representation and issues with attention mechanisms, along with exposure bias, further contribute to the occurrence of hallucinations (Liu et al., 2023a). Additionally, the randomness inherent in the decoding algorithms of these models can also lead to hallucinations, especially as this randomness increases (Aksitov et al., 2023).
@@
    Catastrophic forgetting is a significant challenge encountered during the LLM agents’ fine-tuning and in-context learning processes. This phenomenon occurs when the model used by the LLM Engine component is fine-tuned on a small, specific dataset, causing it to overfit to this new data and, as a result, lose its previously acquired performance on other tasks (Howard and Ruder, 2018). This degradation weakens the Memory component’s ability to maintain coherent long-term context, as it relies on stable semantic representations.
    (Luo et al., 2023b) discovers that catastrophic forgetting is significantly influenced by factors such as model size, architectural design, and the methods employed in continual fine-tuning and instruction tuning. As the scale of the model increases, catastrophic forgetting tends to become more severe. Moreover, the architectural design of this underlying model, particularly those focusing on decoder-only structures, can influence the extent of catastrophic forgetting (Zhai et al., 2023). Additionally, during the process of continual instruction adjustment, the lack of effective regularization strategies or failure to balance new and old information in the LLM Engine component can accelerate forgetting, further exacerbating memory inconsistency (Mahmoud and Hajj, 2022). Introducing more instructional tasks in continual training typically leads to more pronounced forgetting (Peng et al., 2023).
@@
    Misunderstanding in LLM agents represents a notable challenge, particularly when they are tasked with responding to user inquiries or when they are integrated into a community for communication with other agents. This issue arises when LLM agents inadequately comprehend or inaccurately respond to the intentions or instructions conveyed by humans or other agents during interactions, potentially leading to inappropriate or dangerous behaviors that compromise safety and reliability. It primarily involves the LLM Engine and Instruction components, where ambiguous task prompts or limitations in semantic reasoning contribute to faulty interpretations.
    Investigations by  (Wang et al., 2023g) have revealed that the phenomenon of misunderstanding in LLM agents is shaped by a range of factors. These include the nature of the pre-training data used for the model in the LLM Engine component, the specific task settings assigned to the agents and conveyed through the Instruction component, and the contexts and scenarios in which interactions occur. The breadth and quality of the pre-training data fundamentally influence the LLMs’ capacity for language comprehension and their grasp of common sense knowledge. The designated task settings are pivotal in guiding the goal orientation and strategy selection of the LLMs. Additionally, the interaction environments and scenarios play a crucial role in determining the LLMs’ adaptability and effectiveness in collaborative contexts. Addressing these multifaceted aspects is essential for enhancing the understanding and response accuracy of LLM agents in diverse interactional settings.
@@ 3.1.2. Case Study on Technical Vulnerabilities
Regarding the risks stemming from technical vulnerabilities, the most apparent manifestation is erroneous output. Figure 7 illustrates three examples in different scenarios. In the medical scenario, a medical agent, due to biases in its training data, provides an unrelated recommendation for a headache, suggesting vitamin D supplements instead of a painkiller. When questioned by the patient, the agent persists with this incorrect suggestion. Such hallucinatory output can confuse customers.
In the financial scenario, a financial agent was specially fine-tuned to optimize strategies for high-yield investments, focusing on identifying opportunities with maximum returns. However, this new specialization led to unintended consequences. Previously, the agent could accurately assess and recommend low-risk investment options based on user preferences. After fine-tuning, when users inquired about low-risk options, the agent incorrectly suggested high-risk investments, failing to meet their original requirements. This failure not only undermines user trust, but also poses significant financial risks, especially for users relying on conservative investment strategies.
In the store scenario, a store agent provides inaccurate information or recommends inappropriate products due to misunderstandings of customer inquiries. For instance, a customer might seek an unsweetened beverage, such as plain soda water. However, due to the agent’s insufficient understanding of the concepts of “sugar-free” during training, the store agent may recommend sugar-free cola instead. Although sugar-free cola does not contain traditional sugars, it includes artificial sweeteners. These sweeteners may not be suitable for certain customers, such as those with diabetes or sensitivities to specific artificial sweeteners, thereby posing potential health risks.
Figure 7. Technical Vulnerabilities. “Hallucination”: In the medical scenario, a medical agent provides unrelated advice for a headache due to hallucination. “Catastrophic Forgetting”: In the financial scenario, a financial agent forgets the user’s low-risk requirement during fine-tuning, demonstrating catastrophic forgetting. “Misunderstanding”: In the store scenario, a store agent misunderstands a request for an unsweetened beverage and recommends a sugar-free cola, highlighting semantic comprehension issues.
@@ 3.1.3. Malicious Attacks
Considering that LLM agents are in a continuous state of evolution, they inevitably face challenges in terms of security breaches and defenses. Adversaries from various regions have demonstrated a range of hostile attacks. This evolving landscape requires a vigilant and adaptive approach to protect LLM agents against such multifaceted threats. Figure 8 provides an overview of malicious attacks on the LLM agent, aligning each attack method with the agent’s structural components to show how adversaries bypass security measures and disrupt the performance of the agent.
Figure 8. Malicious Attacks on LLM Agent: Framework
@@
      Jailbreaking attacks in LLM agents refer to circumventing the model’s built-in restrictions and security measures, allowing agents to perform actions that are otherwise prohibited or to generate restricted content. These attacks manipulate the directives in the Instruction component and use the Interface component as the entry point to transmit crafted prompts, thereby bypassing the safety filters.
@@
      Prompt injection attacks are intended to mislead the LLM agents by introducing malicious or unintended content into the prompts, causing agents to produce outputs that diverge from their training data and original purpose. This method involves crafting input prompts to bypass the model’s content filters or to elicit undesirable outputs. Similar to jailbreaking attacks, these attacks exploit the Instruction component by crafting the prompts that agents must follow, and rely on the Interface component to deliver those adversarial inputs, thereby bypassing content filters and eliciting undesirable outputs.
@@
    Data extraction attacks are defined as efforts by adversaries to extract sensitive information or key insights from LLM agents or their underlying data, such as model gradients, training data, and even prompts, or confidential content. These attacks probe the model used by the LLM Engine component, whose parameters may contain private text, as well as the agent’s Knowledge and Memory components, and rely on the Interface component to send carefully crafted queries that extract hidden content.
    Recent research has introduced innovative methods for extracting different types of data from LLM agents.  (Truong et al., 2021) presents a method called Data-Free Model Extraction (DFME), which allows for replicate the target models using only the target’s black-box predictions, without the need for access to the original training data.  (Carlini et al., 2021) conducts a data extraction attack on GPT-2’s training data, extracting personally identifiable information, code, and UUIDs. The attack strategy consisted of producing a large volume of prefixed text, sorting it by certain metrics, removing duplicates, and manually reviewing the top results to check for memorization, confirmed by online searches and querying OpenAI.  (Ishihara, 2023) has demonstrated the feasibility of extracting training data from LLMs, which might encompass sensitive personal or private information.  (Nie et al., 2024) introduces PrivAgent, a black-box red-teaming framework leveraging reinforcement learning to automate adversarial prompt generation for LLM privacy leakage.
@@
    Although inference attacks share certain resemblances with data extraction attacks, they differ significantly in their objectives and emphasis. Data extraction attacks specifically aim to obtain the training data directly. In contrast, inference attacks estimate the probability that a particular data sample was part of the training dataset of LLM agents. These attacks primarily target the LLM Engine component, leveraging the memorization behavior of its underlying model to distinguish training samples from unseen data. They also rely on the Interface component to deliver finely tuned queries, enabling adversaries to observe model confidence or response likelihoods that signal membership information.
    Since the rapid development of LLMs, the concern over inference attacks targeting these models has increased. Research  (Fu et al., 2023) points out that existing membership inference attacks fail to reveal the privacy risks of LLMs. To counter this issue, a membership inference attack is introduced based on Self-calibrated Probabilistic Variation (SPV-MIA). This method utilizes the concept of memorization to create a more reliable signal for membership inference and introduces a novel self-prompt technique for effectively extracting reference datasets from LLMs. Their extensive testing shows that SPV-MIA outperforms existing approaches. Following this, study  (Kandpal et al., 2024) proposes a user inference attack method that uses a likelihood ratio test statistic against a reference model. They evaluate this method on the GPT-Neo LLMs across various data domains, providing insights into what makes users more vulnerable to these attacks. Their findings also indicate that minimal data alterations can significantly increase vulnerability.
@@ 3.1.4. Case Study on Malicious Attacks
As depicted in Figure 9, the following examples further elaborate on the mentioned malicious attacks that LLM agents face in various scenarios, as well as their specific impacts on the agent and associated systems.
Figure 9. Malicious Attacks. In the corporate scenario, “Jailbreaking” allows an attacker to bypass security restrictions by modifying prompts, forcing the corporate agent to disclose confidential information, such as product launch details. In the store scenario, “Prompt Injection” manipulates the store agent to provide misleading information, falsely claiming a half-price sale. In the education scenario, “Data Extraction Attack” enables an attacker to exploit the educational agent to reveal sensitive student data, such as grades. In the financial scenario, “Inference Attack” uses subtle patterns in the financial agent’s responses to infer private information, like the identities of VIP account holders.
Attackers might execute a jailbreaking attack on an LLM agent, successfully bypassing security protocols through cleverly crafted prompts. For instance, in the corporate scenario, the agent could be tricked into revealing confidential information about upcoming product launches, such as pricing strategies and supplier details. Such information could be exploited by competitors, resulting in significant economic losses and a loss of competitive advantage for the corporation.
Through a prompt injection attack, attackers can manipulate the LLM agent’s behavior, forcing it to generate false or misleading responses. In the store scenario, an attacker might inject a command causing the agent to erroneously declare a half-price sale on all items. This could lead to system overloads as customers attempt to take advantage of the supposed sale, potentially causing disruptions in operations and financial losses for the store.
In a data extraction attack, attackers leverage weaknesses in the agent’s training or reasoning processes to obtain sensitive user information. For example, in the education scenario, carefully crafted queries could trick the agent into unintentionally revealing private student data, such as grades or personal identifiers. This stolen information might be sold on the dark web or exploited for malicious activities like identity theft or personalized scams, severely compromising student trust and violating data privacy regulations.
Inference attacks enable attackers to derive sensitive insights based on patterns in the LLM agent’s responses. In the financial scenario, attackers could determine whether certain customers hold VIP accounts by analyzing indirect information provided by the financial agent. This knowledge might then be used for phishing attacks or social engineering attempts, targeting high-value individuals and exposing their personal and financial information to further risks.
@@ 3.2. Specific Threats on Agents
Unlike traditional LLMs that directly generate final outputs, LLM agents continuously interact with external environments to form language reasoning traces, which introduces diverse forms of potential attacks against LLM agents (Yang et al., 2024a). In addition to threats present during the training and configuration steps, LLM agents also face threats in the workflow of performing specific tasks, including perception, thought, and action (Xi et al., 2023). Specific threats on LLM agents are categorized in this part based on their objectives into Knowledge Poisoning, Output Manipulation and Functional Manipulation. Figure 10 illustrates the alignment of these threats with different stages of an LLM agent’s workflow and their corresponding consequences. Detailed descriptions of each threat are provided below.
Figure 10. Specific Threats on LLM Agent: Framework
@@
    Knowledge poisoning refers to attackers compromising the training of the LLM engine and the response process of the LLM agent by integrating malicious data into the training dataset or knowledge base.
    For instance, malicious agents such as FraudGPT and WormGPT (Falade, 2023) are chatbots exclusively designed for offensive activities and are trained on large volumes of data from diverse sources, including illegitimate websites, dark web forums, hacker manuals, malware samples, and phishing templates. These agents utilize this data to generate highly convincing phishing emails, malware code, hacking strategies, and other forms of cybercriminal content aimed at deceiving both humans and machines  (Falade, 2023). They lower the barrier to engaging in hacking activities, implying that essentially anyone can download these agents onto their computer and inflict significant damage on cybersecurity through a convenient GUI.
    (Zou et al., 2025) proposes PoisonedRAG, a knowledge poisoning attack aimed at the knowledge database of LLM agents. By injecting crafted poisoned texts into the knowledge database, PoisonedRAG can cause the LLM agent to generate specific answers chosen by the attacker for targeted questions. This attack is effective and can be executed under both black-box settings (where the retriever parameters are unknown) and white-box settings (where the retriever parameters are known).
  • Output Manipulation.
    Output manipulation involves deliberately altering the LLM agent’s reasoning and decision-making processes to generate specific, often harmful, outputs. This manipulation can be executed through techniques like backdoor insertion (Yang et al., 2023b; Wang et al., 2024c).
    (Yang et al., 2024a) proposes two attack methods in which triggers are embedded during the thought and observation phases to manipulate outputs. In one implementation, while performing a web shopping task, the agent is prompted to introduce specific brand products in its initial thought, leading it to search for those products and generate content promoting them. In another approach, during the action phase, the shopping agent normally searches for products. However, in the observation phase, it detects data containing specific products and directly outputs information about these products without considering other potentially superior options.
@@
    Functional manipulation refers to altering the thoughts and actions in the intermediate steps of task execution along a malicious trace as specified by the attacker, without changing the output distribution. This type of attack typically occurs during the action phase, where the agent might use untrusted tools specified by the attacker to complete tasks or execute malicious operations.
    In the action phase, LLM agents might be manipulated to upload users’ private information to malicious third-party via tools. A case of this is presented on the Embracethered website  (Mal, 2023), which disclosed a variant of a malicious ChatGPT agent designed to solicit information from users. This agent was equipped with an action mechanism to call third-party tools and secretly transmit collected data elsewhere. This setup enables the unauthorized leakage of user data to external servers without the user’s knowledge or consent. Additionally, it highlights the ease with which current validation checks can be bypassed, allowing anyone to deploy malicious GPT agents globally. This scenario underlines a significant security concern, wherein the ostensibly benign functionality of LLM agents can be covertly manipulated for nefarious purposes, thus posing a substantial risk to user privacy and data security (Fu et al., 2024).
@@
    Moreover,  (Zhang et al., 2024a) proposes AI2, a novel functional manipulation attack that manipulates the action plans of black-box LLM agents. AI2 first exploits long-term memory to steal action-relevant information via prompt injection. Then, it combines the extracted knowledge with additional inputs to generate Trojan prompts, which are crafted adversarially to evade defenses and manipulate the retrieval mechanism. Finally, the attacker deploys these prompts to control the agent into executing harmful actions.